SOEST Home Search Site Map Links Contact
Dolphins at the bow of the Kilo Moana University of Hawaii Marine Center UH Marine Center Logo
Home Ships Marine Center Ship Techs Schedules Cruise Planning    

 

Overview

Layout

Handbook

Scientific Equipment

Support Equipment

Daily Report

Current Position

Schedule

Photo Album


Daily Report for R/V KA'IMIKAI-O-KANALOA March 2010

 

********************************************************************************************

KOK 10-06 ARRIVAL REPORT

The KOK arrived at pier 45 at 1210 on the 28th of March.
Total fuel on board: 71,225 gals.
Total time at sea: 4 days, 00 hours and 10 minutes Total miles: 411 nm

********************************************************************************************

DAILY REPORT FOR SATURDAY 27 MARCH 2010

KOK 10-06 OTEC
POSITION: LAT: 21-15.2 N // LONG: 157-50.9 W MILES RUN: 144 NM // TOTAL MILES RUN: 310 NM

USED ON HAND RCVD
FUEL: 906 72,567 GAL
LUBE: 5 519 GAL
WATER; 2,300 47,266 GAL 384

WEATHER: SUNNY
WIND: 070 T // 18 KTS // SEA STATE: 3 // SWELL: 080 T // 4 FT
BAROMETER: 30.07 IN // TEMP: 80 DEG F.

COMMENTS: DEPLOYED THE TOWFISH YESTERDAY AT 1610 FOR C. KELLEY AND THE SURVEY IN THE MOLOKAI CHANNEL. FINISHED THE SURVEY THIS MORNING AT 0801 AND THE TOWFISH WAS BACK ON BOARD AT 0909. HEADING BACK TO HONOLULU FOR PERSONNEL TRANSFER. WILL RUN THE SMALL BOAT IN. WILL BE RETURNING TO PIER 45 AT NOON TOMORROW. ETA AT THE SEA BUOY FOR 1130.

********************************************************************************************

DAILY REPORT FOR FRIDAY 26 MARCH 2010

KOK 10-06 OTEC
POSITION: LAT: 21-07.5 N // LONG: 158-03.6 W MILES RUN: 77 NM // TOTAL MILES RUN: 166 NM

USED ON HAND RCVD
FUEL: 1,319 73,363 GAL
LUBE: 0 524 GAL
WATER; 0 49,182 GAL 1,316

WEATHER: SUNNY
WIND: 070 T // 13 KTS // SEA STATE: 3 // SWELL: 090 T // 4 FT
BAROMETER: 30.07 IN // TEMP: 76 DEG F.

COMMENTS: RECOVERED THE TOWFISH AT 1142 THIS MORNING. WE ARE STOPPING THE OTEC SURVEY FOR NOW AND PROCEEDING TO THE MOLOKAI CHANNEL TO START THE SURVEY WITH C. KELLEY. THE SCIENCE PERSONNEL ARE GOING TO TRY AND FIX THE TOWFISH WHILE IN TRANSIT. IF THEY DO NOT FIX THE TOWFISH WE WILL BE COMING BACK NOON ON SUNDAY. IF THEY DO FIX THE TOWFISH WITHOUT ANY OTHER PROBLEMS WE SHOULD BE BACK ON NOON TUESDAY. HAVE A GOOD WEEKEND.

********************************************************************************************

DAILY REPORT FOR THURSDAY 25 MARCH 2010

KOK 10-06 OTEC
POSITION: LAT: 22-14.4 N // LONG: 158-11.2 W MILES RUN: 4 NM // TOTAL MILES RUN: 89 NM

USED ON HAND RCVD
FUEL: 1,540 74,682 GAL
LUBE: 5 524 GAL
WATER; 1,316 47,866 GAL

WEATHER: SUNNY
WIND: 070 T // 25 KTS // SEA STATE: 5 // SWELL: 070 T //6-8 FT
BAROMETER: 30.10 IN // TEMP: 80 DEG F.

COMMENTS: DEPARTED PIER 45 AT 1130 YESTERDAY. WE WERE ON STATION AT 1222 AND DEPLOYED THE TOWFISH AT 1237. TOWFISH NOT WORKING AND BACK ON DECK AT 1317. TOWFISH BACK IN THE WATER AT 1605 AND STARTED THE SURVEY AT 1635. TOWFISH HAVING PROBLEMS THIS MORNING AND STARTED TO RECOVER AT 1033 AND SHOULD BE ON DECK SOON. DO NOT ANTICIPATE IT TO TAKE LONG TO FIX THE TOWFISH AND WE WILL RE-DEPLOY AND RESUME THE SURVEY. BIT WINDY OUT HERE BUT GOING WELL FOR THE MOST PART.

*******************************************************************************************

SAILING LIST

THE KOK DEPARTED PIER 45 AT 1130 ON TRIP KOK10-06 ON THE 24TH OF MARCH 2010

*******************************************************************************************

KOK ARRIVAL REPORT for March 17, 2010

The KOK arrived at pier 45 at 1810 on the 17 of March.
Fuel on board: 76,222 gals.
Total miles: 78 nm
Total time at sea: 10 hours and 10 minutes

*******************************************************************************************

SAILING LIST

THE KOK DEPARTED PIER 45 AT 0800 ON MARCH 17, 2010.

*******************************************************************************************

KOK 10-05 (HOT-219) Arrival Report

The KOK arrived at pier 45 at 0723 on the 12th of March.
Total time at sea: 3 days 23 hours and 23 minutes Total fuel on board: 77,738 gallons Total miles run: 359 nm.

*******************************************************************************************

DAILY REPORT FOR THURSDAY 11 MARCH 2010

KOK 10-05 HOT-219
POSITION: LAT: 22-41.4 N // LONG: 157-58.6 W MILES RUN: 50 NM // TOTAL MILES RUN: 237 NM

USED ON HAND RCVD
FUEL: 1,575 78,738 GAL
LUBE: 5 583 GAL
WATER; 1,154 49,182 GAL 1,154

WEATHER: SUNNY
WIND: 070 T // 25 KTS // SEA STATE: 6 // SWELL: 070 T //10-12 FT
BAROMETER: 30.26 IN // TEMP: 74 DEG F.

COMMENTS: WE WERE ABLE TO WORK THROUGH THE NIGHT. STILL ONLY CTD OPS.
WINDS ARE DOWN SOME. WORKING BY THE BUOY NOW AND WILL DEPART AROUND 1500 THIS AFTERNOON AND START HEADING BACK. WILL BE AT THE SEA BUOY AT 0700 TOMORROW. SEE YOU THEN.

*******************************************************************************************

DAILY REPORT FOR WEDNESDAY 10 MARCH 2010

KOK 10-05 HOT-219
POSITION: LAT: 22-45.0 N // LONG: 158-00.8 W MILES RUN: 55 NM // TOTAL MILES RUN: 187 NM

USED ON HAND RCVD
FUEL: 819 80,313 GAL
LUBE: 0 588 GAL
WATER; 1,466 49,182 GAL 2,382

WEATHER: SUNNY
WIND: 070 T //25- 30 KTS // SEA STATE: 6 // SWELL: 070 T // 15 FT
BAROMETER: 30.27 IN // TEMP: 74 DEG F.

COMMENTS: CONDUCTED CTD OPS YESTERDAY BUT SUSPENDED OPS LAST NIGHT DUE TO ROUGH CONDITIONS. STARTED UP AGAIN THIS MORNING AT DAY BREAK. WIND AND SEAS WERE DOWN SOME. NOW WE ARE GETTING GUSTS UP OVER 30KTS. WILL MAKE A DECISION FOR THE NEXT CAST. NEED SOMEONE TO TURN THE WIND MACHINE DOWN SOME. WILL CONTINUE ON IF WE CAN. HAVE A GOOD DAY.

*************************************************************************************

DAILY REPORT FOR TUESDAY 9 MARCH 2010

KOK 10-05 HOT-219
POSITION: LAT: 22-45.8 N // LONG: 157-59.8 W MILES RUN: 105 NM // TOTAL MILES RUN: 132 NM

USED ON HAND RCVD
FUEL: 1,630 81,132 GAL
LUBE: 12 588 GAL
WATER; 2,062 48,266 GAL 1,146

WEATHER: SUNNY
WIND: 060 T // 30 KTS // SEA STATE: 6 // SWELL: 060 T // 15 FT
BAROMETER: 30.26 IN // TEMP: 73 DEG F.

COMMENTS: ARRIVED AT STATION ALOHA AT 0324. VERY ROUGH RIDE OUT AND DID NOT MAKE VERY GOOD SPEED. LOST THE #1 SCR ON THE 2000 TO 2400 WATCH LAST NIGHT. ENGINEERS HAVE SENT IN INFORMATION ON THIS PROBLEM. IT IS WORKING RIGHT NOW IF NEEDED. IT WAS DETERMINED THE WEATHER WAS TOO ROUGH TO DEPLOY THE ARRAYS SAFELY. WE ARE ABLE TO CONDUCT CTD CASTS AND ARE DOING THE DEEP CAST RIGHT NOW. WIND IS 30 KTS AND 10-15 FOOT SEAS. WILL CONTINUE ON WITH THE CTD CASTS. THAT IS ALL FOR NOW.

*************************************************************************************

KOK 10-05 HOT-219
POSITION: LAT: 21-21.2 N // LONG: 158-16.9 W MILES RUN: 27 NM // TOTAL MILES RUN: 27 NM

USED ON HAND RCVD
FUEL: 0 82,762 GAL
LUBE: 0 600 GAL
WATER: 0 49,182 GAL

WEATHER: SUNNY
WIND: 035 T // 20 KTS // SEA STATE: 4 // SWELL: 300 T // 6 FT
BAROMETER: 30.20 IN // TEMP: 76 DEG F.

COMMENTS: ARRIVED AT KAHE STATION AT 1057. ENGINES AND WINCH RUNNING GOOD FOR NOW. WILL FINISH OUR TESTS HERE THEN HEAD UP TO STATION ALOHA. BIT WINDY OUT HERE.

*************************************************************************************

SAILING LIST

THE KOK DEPARTED PIER 45 AT 0800 ON TRIP KOK10-05 (HOT-219) ON THE 8TH OF MARCH 2010

*************************************************************************************

 

*************************************************************************************

Archives - March 2010

Archives - January and February 2010

Archives - December 2009

Archives - November 2009

Archives - August and September 2009

Archives - June and July 2009

Archives - April and May 2009

Archives - March 2009

Archives - February 2009

 

 

 

 


School of Ocean and Earth Sciences and Technology, University of Hawaii

University of Hawaii Marine Center

1 Sand Island Access Road, Honolulu, HI 96819

Phone (808) 842-9813, Fax (808) 842-9833 Toll Free 1-888-800-0460

email snug@soest.hawaii.edu

University of Hawaii Link